From the final answer when I multiply 2 by 14.2 and divide 5000 by 28.4, my answer is 176 days
Answer #1

2*14.2= 28.4

W(gett after multiplication)

Now 5000/ 28.4= 176.056  

Is the required ans.
Answer #2

Solution

Given Operations:

  1. Multiply 2 × 14.2:

    2×14.2=28.4

  2. Divide 5000 ÷ 28.4:

    5000÷28.4176.056 (rounded to 176 days)

Final Answer:
The result is 176 days


answered by: anonymous